OpenStudy (anonymous):

The area of a triangle is found using the formula A = 1/2 bh, where b is the base and h is the height of the triangle. Solve this formula for the height h. Then find the height of a triangular flag if the Area = 63mm2 and the base is 9 mm.

OpenStudy (anonymous):

A = 1/2 bh 63mm^2 = 1/2 (9mm) h 63mm^2 = 4.5mm (h) divide both sides by 4.55mm so you would get... 14mm = h check: A = 1/2 (9mm)(14mm) A = 63mm^2